# Conditional Distribution Compression via the Kernel Conditional Mean Embedding

### v3

@@ 1 Introduction
No existing distribution compression method targets the family of conditional distributions $\mathbb{P}_{Y\mid X}$ of a labelled dataset. Intuitively, we posit that directly compressing the conditional distribution should be preferable to indirect joint compression, much as direct conditional density estimation outperforms approaches based on separate joint and marginal estimates [17]. To address the gap, we require the kernel conditional mean embedding (KCME), denoted $\mathcal{\mu}_{Y\mid X}$. The KCME provides a way to embed the family of conditional distributions $\mathbb{P}_{Y\mid X}$ into an RKHS, and is a widely used technique for non-parametric modelling of complex conditional distributions. Given $n$ labelled samples, it can be consistently estimated with a computational cost of $\mathcal{O}(n^{3})$ [18, 19]. Despite this high cost, the KCME has been successfully applied in various fields, including conditional distribution testing [19, 20], conditional independence testing [19, 21], conditional density estimation [22, 23, 24], likelihood-free inference [25], Bayesian optimisation [26], probabilistic inference [18, 23, 27, 28, 29], calibration of neural networks [30], reinforcement learning [31, 32, 33], and as a consistent multi-class classifier [34]. Proofs of all our theoretical results are in Section B.
@@ 2 Preliminaries
The KCME can be written as the composition of a function $F_{Y\mid X}:\mathcal{X}\to\mathcal{H}_{l}$ and the random variable $X:\Omega\to\mathcal{X}$, i.e. $\mu_{Y\mid X}=F_{Y\mid X}\circ X$ (Theorem 4.1, [19]). For consistency in notation, throughout the remainder of this work, whenever we refer to the KCME $\mu_{Y\mid X}$, we mean $F_{Y\mid X}$. The KCME can be estimated directly [19, 31] using i.i.d. samples from the joint distribution $\mathcal{D}$ as
@@ 3.1 Joint Kernel Herding
By inducing an RKHS $\mathcal{H}_{k}\otimes\mathcal{H}_{l}$ with the tensor product kernel $k(\cdot,\cdot)l(\cdot,\cdot)$, one can modify the Kernel Herding [3] algorithm to instead target the joint distribution. Assuming we are at the $(m+1)^{\text{th}}$ iteration, having already constructed a compressed set of size $m$, $\mathcal{C}=\{(\tilde{\bm{x}}_{j},\tilde{\bm{y}}_{j})\}_{j=1}^{m}$, the next pair is chosen as the solution to the optimisation problem
@@ 3.2 Joint Kernel Inducing Points
Defining $\bm{X}:=[\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{n}]^{T}$, $\bm{Y}:=[\bm{y}_{1},\bm{y}_{2},\dots,\bm{y}_{n}]^{T}$, $\tilde{\bm{X}}:=[\tilde{\bm{x}}_{1},\tilde{\bm{x}}_{2},\dots,\tilde{\bm{x}}_{m}]^{T}$, and $\tilde{\bm{Y}}:=[\tilde{\bm{y}}_{1},\tilde{\bm{y}}_{2},\dots,\tilde{\bm{y}}_{m}]^{T}$, we target the $\text{JMMD}(\mathbb{P}_{X,Y},\tilde{\mathbb{P}}_{X,Y})$ by solving the optimisation problem
@@
[…]
@@ 4.1 The Average Maximum Conditional Mean Discrepancy
In the above definition, the expectation is taken with respect to a distinct probability measure $\mathbb{P}_{X^{*}}\neq\mathbb{P}_{X},\mathbb{P}_{X\prime}$, which broadens its applicability compared to the KCD. See Section C.2 for an illustrative example.
@@ 4.2 Average Conditional Kernel Herding
Assuming we are at the $m^{\text{th}}$ iteration, having already constructed the compressed set $\mathcal{C}=\{(\tilde{\bm{x}}_{j},\tilde{\bm{y}}_{j})\}_{j=1}^{m-1}$, we expand the compressed set $\mathcal{C}=\mathcal{C}\cup(\bm{x},\bm{y})$ by solving
$\underset{(\bm{x},\bm{y})\in\mathcal{X}\times\mathcal{Y}}{\arg\min}\;\;\left\{\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\|\tilde{\mu}^{\mathcal{C}}_{Y\mid X=\bm{x}^{\prime}}\right\|^{2}_{\mathcal{H}_{l}}\right]-2\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\langle\mu_{Y\mid X=\bm{x}^{\prime}},\tilde{\mu}^{\mathcal{C}}_{Y\mid X=\bm{x}^{\prime}}\right\rangle_{\mathcal{H}_{l}}\right]\right\}.$ (8)
@@
Letting $\mathcal{G}^{\mathcal{D}}(\bm{x},\bm{y})$ be the estimate of the objective in (8) computed using $\mathcal{D}$, we obtain
$\mathcal{G}^{\mathcal{D}}(\bm{x},\bm{y})$ $=\frac{1}{n}\text{Tr}(\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T})-\frac{2}{n}\text{Tr}(\bar{K}(\bm{x})\tilde{W}(\bm{x})\bar{L}(\bm{y})^{T}),$ (9)
where we define $\bar{K}(\bm{x})\in\mathbb{R}^{n\times m}$ to have entries $k(\bm{x}_{i},\tilde{\bm{x}}_{j})$, where $\bm{x}_{i}\in\mathcal{D}$ and $\tilde{\bm{x}}_{j}\in\mathcal{C}$, for $i=1,\dots,n$ and $j=1,\dots,m$. Similarly, $\tilde{K}(\bm{x})\in\mathbb{R}^{m\times m}$ is defined to have entries $k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})$ for $\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j}\in\mathcal{C}$, with $i,j=1,\dots,m$. We then set $\tilde{W}(\bm{x}):=(\tilde{K}(\bm{x})+\lambda I)^{-1}$. Analogous definitions hold for the response kernel $l$, yielding matrices $\bar{L}(\bm{y})\in\mathbb{R}^{n\times m}$ and $\tilde{L}(\bm{y})\in\mathbb{R}^{m\times m}$. To construct a compressed set of size $m$, ACKH can be shown to have an overall time complexity of $\mathcal{O}(m^{4}+m^{3}n)$; see Section D.2.3.
@@ 4.3 Average Conditional Kernel Inducing Points
Unlike JKH, where updates depend only on the newest pair in the compressed set, the presence of an inverse in the objective (9) prevents a similar simplification.  [^1] Through the method of bordering [43], it is technically possible to update the inverse of a growing matrix, minimising re-computation. Unfortunately, bordering is a highly numerically unstable procedure, and kernel matrices are often very ill-conditioned in practice. This leads to a quartic dependence of ACKH on the compressed set size $m$, which is a fairly significant limitation. By instead optimising each pair in the compressed set at once, we can reduce this to just cubic dependence on $m$. We refer to this algorithm as Average Conditional Kernel Inducing Points (ACKIP), solving the optimisation problem
$\underset{(\tilde{\bm{X}},\tilde{\bm{Y}})\subset\mathcal{X}\times\mathcal{Y}}{\arg\min}\;\;\left\{\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\|\tilde{\mu}^{(\tilde{\bm{X}},\tilde{\bm{Y}})}_{Y\mid X=\bm{x}^{\prime}}\right\|^{2}_{\mathcal{H}_{l}}\right]-2\mathbb{E}_{(\bm{x}^{\prime},\bm{y}^{\prime})\sim\mathbb{P}_{X,Y}}\left[\tilde{\mu}^{(\tilde{\bm{X}},\tilde{\bm{Y}})}_{Y\mid X=\bm{x}^{\prime}}(\bm{y}^{\prime})\right]\right\}$ (10)
@@ 5 Experiments
Building on the experimental setup of Kernel Herding [3], we demonstrate how the methods proposed in this paper can be applied to compress the true conditional distribution (under certain conditions on the distribution and kernels) and the empirical conditional distribution. Specifically, we report the root mean square error $\text{RMSE}(\mathcal{C}):=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(\mathbb{E}[h(Y)\mid X=\bm{x}_{i}]-\langle\hat{\mu}^{\mathcal{C}}_{Y\mid X=\bm{x}_{i}},h\rangle_{\mathcal{H}_{l}}\right)^{2}}$, across a range of test functions $h:\mathcal{Y}\to\mathbb{R}$. This aligns with the standard applications of the KCME [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], where one estimates the conditional expectation of a function of interest $h$. When the exact value of the conditional expectation is unavailable, we approximate $\mathbb{E}[h(Y)\mid X=\bm{x}_{i}]$ via its full-data estimate, $\langle\hat{\mu}^{\mathcal{D}}_{Y\mid X=\bm{x}_{i}},h\rangle_{\mathcal{H}_{l}}$. Note that when $h$ is the identity function, the estimate reduces to the familiar regression setting, and when $h$ is an indicator function, it corresponds to classification. For full details of the experiments, including results on additional test functions omitted from the main text due to space constraints, see Section C.
@@ 5.2.1 Continuous Conditional Distributions
Synthetic: We design a challenging dataset with a highly non-linear feature-response relationship and pronounced heteroscedastic noise, referring to it as Heteroscedastic going forward. Let $\mathbb{P}_{X}=\mathcal{N}(0,3^{2})$ and $\mathbb{P}_{Y\mid X=\bm{x}}=\mathcal{N}(f(\bm{x}),\sigma^{2}(\bm{x}))$, with $f(\bm{x}):=\sum_{i=1}^{4}a_{i}\exp\left(-\frac{1}{b_{i}}(\bm{x}-c_{i})^{2}\right)$, and $\sigma^{2}(\bm{x}):=\sigma_{1}^{2}+\left|\sigma_{2}^{2}\sin(\bm{x})\right|$. Figure 1 compares a compressed set constructed by ACKIP, with the random sample which initialised the optimisation procedure. It demonstrates how random sampling fails to adequately represent key areas of the data cloud, and how ACKIP can construct a better representation. In Figures 6 and 7 we see that ACKIP attains the lowest RMSE across three of the four test functions. For the remaining function, all methods exhibit relatively similar performance. We also note that JKIP outperforms or achieves similar performance versus JKH across the test functions.
Figure 6: RMSE versus size of compressed set for the Heteroscedastic data with parameters set as $\bm{a}:=[3,-3,6,-6]^{T}$, $\bm{b}:=[1,0.1,2,0.5]^{T}$, $\bm{c}:=[-5,-2,2,5]^{T}$, $\sigma_{1}^{2}=0.1$ and $\sigma_{2}^{2}=0.75$. The RMSE is calculated against the true value of the conditional expectations: the performance of the full data (purple) is hence also highlighted here.
@@ 5.2.2 Discrete Conditional Distributions
Conditional distributions may also be discrete, as in classification settings where responses take one of $C$ distinct values $\bm{y}\in\{0,1,\dots,C\}$. In such cases, applying an indicator kernel on the response space enables the KCME to serve as a consistent multi-class classifier, in contrast to methods like SVCs and GPCs [34]. The use of the indicator kernel renders standard gradient-based optimisation inapplicable on the response space, necessitating an alternative optimisation strategy (see Section C.1.3 for details). Figure 8 illustrates the strong performance of ACKIP on a challenging, synthetic, imbalanced four-class classification dataset, which we refer to as Imbalanced. For full details see Section C.1.3. We see that the KCME, trained with the compressed set constructed by ACKIP, estimates the class-conditional probabilities with accuracy that very closely matches that of the full dataset at just 3% of the size. In contrast, the figure also exposes the limitations of the herding approach: ACKH performs worse than random on three of the four classes, and JKIP outperforms JKH on three out of the four classes. For additional details and experimental results, including on MNIST, see see Section C.1.3.
@@ A.1.1 Standard Distribution Compression
$\bm{z}_{m+1}=\underset{\bm{z}\in\mathbb{R}^{d}}{\arg\min}\;\;\frac{1}{m+1}\sum_{j=1}^{m}k(\bm{z},\bm{z}_{j})-\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[k(\bm{z},\bm{x}^{\prime})\right],$
after which the compressed set is updated to $\bm{Z}:=[\bm{z}_{1},\bm{z}_{2},\dots,\bm{z}_{m},\bm{z}_{m+1}]^{\top}$, and the process is repeated. This greedy strategy has the advantage of focusing computational effort on optimising one new point at a time while leaving previously selected points fixed. However, it may yield suboptimal solutions, as earlier selections are never revisited or refined. We modify this method to target the joint distribution in Joint Kernel Herding.
Gradient Flow [11] methods address a similar objective by solving a discretised Wasserstein gradient flow of the MMD. In practice, this corresponds to performing gradient descent on all points in the compressed set simultaneously, i.e., solving
$\underset{\bm{Z}\subset\mathbb{R}^{d}}{\arg\min}\;\;\frac{1}{m^{2}}\sum_{i,j=1}^{m}k(\bm{z}_{i},\bm{z}_{j})-\frac{2}{m}\sum_{i=1}^{m}\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[k(\bm{z}_{i},\bm{x}^{\prime})\right].$
Under certain technical convexity assumptions on the objective [11], this method can be shown to converge to the global optimum; in practice, however, one should only expect to obtain a locally optimal solution. We modify this method to target the joint distribution in Joint Kernel Inducing Points.
Kernel Thinning [12, 13, 14] constructs a compressed set that is a proper subset of the original dataset, i.e., all points in the compressed set are drawn directly from the input data. This restriction stems from the method’s original motivation—thinning the output of Markov Chain Monte Carlo (MCMC) methods—where subsampling is commonly referred to as standard thinning. Kernel thinning proceeds via a two-stage procedure targeting the MMD. First, the input dataset is probabilistically split into $2^{m}$ MMD-balanced candidate sets, each of size $\lfloor n/2^{m}\rfloor$, discarding the remaining $n-2^{m}\lfloor n/2^{m}\rfloor$ points if $n$ is not evenly divisible. Second, the best candidate set from this partitioning is selected, and then greedily refined by iteratively replacing points with others from the original dataset whenever doing so improves the MMD. While this construction enables the derivation of convergence rates that are state of the art in the literature, its practical performance has been found to be inferior to alternative approaches; see, for example, Chen et al. [16].
@@
where $\bm{a},\bm{a}^{\prime}\sim\mathbb{P}_{X}$ and $\bm{b},\bm{b}^{\prime}\sim\mathbb{Q}_{X}$. Much like the MMD, the energy distance is zero if and only if $\mathbb{P}_{X}=\mathbb{Q}_{X}$ [10, Theorem 1]. Although Support Points is not initially expressed in kernel form, the energy distance is in fact equivalent to the MMD for a particular choice of negative-definite kernel [48].

### v4

@@ 1 Introduction
No existing distribution compression method targets the family of conditional distributions $\mathbb{P}_{Y\mid X}$ of a labelled dataset. To address the gap, we require the kernel conditional mean embedding (KCME), denoted $\mathcal{\mu}_{Y\mid X}$. The KCME provides a way to embed the family of conditional distributions $\mathbb{P}_{Y\mid X}$ into an RKHS, and is a widely used technique for non-parametric modelling of complex conditional distributions. Given $n$ labelled samples, it can be consistently estimated with a computational cost of $\mathcal{O}(n^{3})$ [72, 60]. Despite this high cost, the KCME has been successfully applied in various fields, including conditional distribution testing [60, 59], conditional independence testing [60, 83], conditional density estimation [65, 73, 68], likelihood-free inference [36], Bayesian optimisation [15], probabilistic inference [72, 73, 69, 71, 70], calibration of neural networks [82], reinforcement learning [29, 57, 8], and as a consistent multi-class classifier [35]. Intuitively, we posit that directly compressing the conditional distribution should be preferable to indirect joint compression, much as direct conditional density estimation outperforms approaches based on separate joint and marginal estimates [40]. Proofs of all our theoretical results are in Section B.
@@ 2 Preliminaries
The KCME can be written as the composition of a deterministic function $F_{Y\mid X}:\mathcal{X}\to\mathcal{H}_{l}$ and the random variable $X:\Omega\to\mathcal{X}$, i.e. $\mu_{Y\mid X}=F_{Y\mid X}\circ X$ (Theorem 4.1, [60]). For consistency in notation, throughout the remainder of this work, whenever we refer to the KCME $\mu_{Y\mid X}$, we mean $F_{Y\mid X}$. The KCME can be estimated directly [60, 29] using i.i.d. samples from the joint distribution $\mathcal{D}$ as
@@ 3.1 Joint Kernel Herding
By inducing an RKHS $\mathcal{H}_{k}\otimes\mathcal{H}_{l}$ with the tensor product kernel $k(\cdot,\cdot)l(\cdot,\cdot)$, one can modify the Kernel Herding [13] algorithm to instead target the joint distribution. First, we assume that $\lVert k(\bm{x},\cdot)l(\bm{y},\cdot)\rVert_{\mathcal{H}_{k}\otimes\mathcal{H}_{l}}=R$ for all $\bm{x}\in\mathcal{X}$ and $\bm{y}\in\mathcal{Y}$, where $R$ is a constant. This condition holds for commonly used stationary kernels such as the Gaussian, Laplace, and Matérn kernels. Then, assuming we are at the $(m+1)^{\text{th}}$ iteration, having already constructed a compressed set of size $m$, $\mathcal{C}=\{(\tilde{\bm{x}}_{j},\tilde{\bm{y}}_{j})\}_{j=1}^{m}$, the next pair is chosen as the solution to the optimisation problem
@@ 3.2 Joint Kernel Inducing Points
We target the $\text{JMMD}(\mathbb{P}_{X,Y},\tilde{\mathbb{P}}_{X,Y})$ by solving the optimisation problem
@@
via gradient descent. In the general case, one cannot compute the expectation above, hence we instead target its empirical alternative, namely $\text{JMMD}(\hat{\mathbb{P}}_{X,Y},\tilde{\mathbb{P}}_{X,Y})$. Defining $\bm{X}:=[\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{n}]^{\top}$, $\bm{Y}:=[\bm{y}_{1},\bm{y}_{2},\dots,\bm{y}_{n}]^{\top}$, $\tilde{\bm{X}}:=[\tilde{\bm{x}}_{1},\tilde{\bm{x}}_{2},\dots,\tilde{\bm{x}}_{m}]^{\top}$, and $\tilde{\bm{Y}}:=[\tilde{\bm{y}}_{1},\tilde{\bm{y}}_{2},\dots,\tilde{\bm{y}}_{m}]^{\top}$, this reduces to targeting
@@ 4.1 The Average Maximum Conditional Mean Discrepancy
In the above definition, the expectation is taken with respect to a distinct probability measure $\mathbb{P}_{X^{*}}\neq\mathbb{P}_{X},\mathbb{P}_{x^{\prime}}$, which broadens its applicability compared to the KCD. See Section C.2 for an illustrative example.
The following theorem establishes that the AMCMD is indeed a proper metric.
@@ 4.2 Average Conditional Kernel Herding
Assuming we are at the $(m+1)^{\text{th}}$ iteration, having already constructed the compressed set $\mathcal{C}=\{(\tilde{\bm{x}}_{j},\tilde{\bm{y}}_{j})\}_{j=1}^{m}$, we expand the compressed set $\mathcal{C}=\mathcal{C}\cup(\bm{x},\bm{y})$ by solving
$\underset{\bm{x},\bm{y}}{\arg\min}\;\;\left\{\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\|\tilde{\mu}^{\mathcal{C}}_{Y\mid X=\bm{x}^{\prime}}\right\|^{2}_{\mathcal{H}_{l}}\right]-2\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\langle\mu_{Y\mid X=\bm{x}^{\prime}},\tilde{\mu}^{\mathcal{C}}_{Y\mid X=\bm{x}^{\prime}}\right\rangle_{\mathcal{H}_{l}}\right]\right\}.$ (8)
@@
Even after applying Lemma 4.7 we still have an expectation in (8) which we may not be able to compute in general. Estimating this expectation using $\mathcal{D}$, we obtain the closed form objective
$\mathcal{G}^{\mathcal{D}}(\bm{x},\bm{y})$ $:=\frac{1}{n}\text{Tr}(\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top})-\frac{2}{n}\text{Tr}(\bar{K}(\bm{x})\tilde{W}(\bm{x})\bar{L}(\bm{y})^{\top}),$ (9)
where we define $\bar{K}(\bm{x})]_{ij}:=k(\bm{x}_{i},\tilde{\bm{x}}_{j})$, $i=1,\dots,n$, $j=1,\dots,m$, $\tilde{K}(\bm{x})]_{ij}:=k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})$, $i,j=1,\dots,m$, and $\tilde{W}(\bm{x}):=(\tilde{K}(\bm{x})+\lambda I)^{-1}$. Analogous definitions hold for the response kernel $l(\cdot,\cdot)$, yielding matrices $\bar{L}(\bm{y})\in\mathbb{R}^{n\times m}$ and $\tilde{L}(\bm{y})\in\mathbb{R}^{m\times m}$. To construct a compressed set of size $m$, ACKH can be shown to have an overall time complexity of $\mathcal{O}(m^{4}+m^{3}n)$; see Section D.2.3.
@@ 4.3 Average Conditional Kernel Inducing Points
Unlike JKH, where updates depend only on the newest pair in the compressed set, the presence of an inverse in the objective (9) prevents a similar simplification.  [^1] Through the method of bordering [21], it is technically possible to update the inverse of a growing matrix, minimising re-computation. Unfortunately, bordering is a highly numerically unstable procedure, and kernel matrices are often very ill-conditioned in practice. This leads to a quartic dependence of ACKH on the compressed set size $m$, which is a fairly significant limitation. By instead optimising each pair in the compressed set simultaneously, we can reduce this to just cubic dependence on $m$.
We refer to this algorithm as Average Conditional Kernel Inducing Points (ACKIP), solving the optimisation problem
$\underset{\mathcal{C}}{\arg\min}\;\;\left\{\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\|\tilde{\mu}^{\mathcal{C}}_{Y\mid X=\bm{x}^{\prime}}\right\|^{2}_{\mathcal{H}_{l}}\right]-2\mathbb{E}_{(\bm{x}^{\prime},\bm{y}^{\prime})\sim\mathbb{P}_{X,Y}}\left[\tilde{\mu}^{\mathcal{C}}_{Y\mid X=\bm{x}^{\prime}}(\bm{y}^{\prime})\right]\right\}$ (10)
@@ 5 Experiments
Building on the experimental setup of Kernel Herding [13], we demonstrate how the methods proposed in this paper can be applied to compress the conditional distribution. We report the root mean square error $\text{RMSE}(\mathcal{C}):=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(\mathbb{E}[h(Y)\mid X=\bm{x}_{i}]-\langle\hat{\mu}^{\mathcal{C}}_{Y\mid X=\bm{x}_{i}},h\rangle_{\mathcal{H}_{l}}\right)^{2}}$, across a range of test functions $h:\mathcal{Y}\to\mathbb{R}$. This aligns with the standard applications of the KCME [72, 60, 59, 83, 65, 73, 68, 36, 15, 69, 71, 70, 82, 29, 57, 8, 35], where one estimates the conditional expectation of a function of interest $h$. When the exact value of the conditional expectation is unavailable, we approximate $\mathbb{E}[h(Y)\mid X=\bm{x}_{i}]$ via its full-data estimate, $\langle\hat{\mu}^{\mathcal{D}}_{Y\mid X=\bm{x}_{i}},h\rangle_{\mathcal{H}_{l}}$. Note that when $h$ is the identity function, the estimate reduces to the familiar regression setting i.e. $\mathbb{E}[Y\mid X]$, and when $h$ is an indicator function, it corresponds to estimating class-conditional probabilities e.g. $\mathbb{P}(Y=0\mid X)$. For full details of the experiments, including results on additional test functions omitted from the main text due to space constraints, see Section C.
@@ 5.2.1 Continuous Conditional Distributions
Synthetic: We design a challenging dataset with a highly non-linear feature-response relationship and pronounced heteroscedastic noise, referring to it as Heteroscedastic going forward. Let $\mathbb{P}_{X}=\mathcal{N}(0,2^{2})$ and $\mathbb{P}_{Y\mid X=\bm{x}}=\mathcal{N}(f(\bm{x}),\sigma^{2}(\bm{x}))$, with $f(\bm{x}):=\sum_{i=1}^{4}a_{i}\exp\left(-\frac{1}{b_{i}}(\bm{x}-c_{i})^{2}\right)$, and $\sigma^{2}(\bm{x}):=\sigma_{1}^{2}+\left|\sigma_{2}^{2}\sin(\bm{x})\right|$. Figure 1 compares a compressed set constructed by ACKIP, with the random sample which initialised the optimisation procedure. It demonstrates how random sampling fails to adequately represent key areas of the data cloud, and how ACKIP can construct a better representation. In Figures 6 and 7 we see that ACKIP attains the lowest RMSE across three of the four test functions. For the remaining function, all methods exhibit relatively similar performance. We also note that JKIP outperforms or achieves similar performance versus JKH across the test functions.
Figure 6: RMSE versus size of compressed set for the Heteroscedastic data with parameters set as $\bm{a}:=[3,-3,6,-6]^{\top}$, $\bm{b}:=[1,0.1,2,0.5]^{\top}$, $\bm{c}:=[-5,-2,2,5]^{\top}$, $\sigma_{1}^{2}=0.1$ and $\sigma_{2}^{2}=0.75$. The RMSE is calculated against the true value of the conditional expectations: the performance of the full data (purple) is hence also highlighted here.
@@ 5.2.2 Discrete Conditional Distributions
Conditional distributions may also be discrete, as in classification settings where responses take one of $C$ distinct values $\bm{y}\in\{0,1,\dots,C\}$. In such cases, applying an indicator kernel on the response space enables the KCME to serve as a consistent multi-class classifier, in contrast to methods like SVCs and GPCs [35]. The use of the indicator kernel renders standard gradient-based optimisation inapplicable on the response space, necessitating an alternative optimisation strategy (see Section C.1.3 for details). Figure 8 illustrates the strong performance of ACKIP on a challenging, synthetic, imbalanced four-class classification dataset, which we refer to as Imbalanced. We see that the KCME, trained with the compressed set constructed by ACKIP, estimates the class-conditional probabilities with accuracy that very closely matches that of the full dataset at just 3% of the size. In contrast, the figure also exposes the limitations of the herding approach: ACKH performs worse than random on three of the four classes, and JKIP outperforms JKH on three out of the four classes. For full details on Imbalanced, and additional experimental results, including on MNIST, see see Section C.1.3.
@@ A.1.1 Standard Distribution Compression
$\bm{z}_{m+1}=\underset{\bm{z}\in\mathbb{R}^{d}}{\arg\min}\;\;\frac{1}{m+1}\sum_{j=1}^{m}k(\bm{z},\bm{z}_{j})-\mathbb{E}_{\bm{x}\sim\mathbb{P}_{X}}\left[k(\bm{z},\bm{x})\right],$
after which the compressed set is updated to $\bm{Z}:=[\bm{z}_{1},\bm{z}_{2},\dots,\bm{z}_{m},\bm{z}_{m+1}]^{\top}$, and the process is repeated. This greedy strategy has the advantage of focusing computational effort on optimising one new point at a time while leaving previously selected points fixed. However, it may yield suboptimal solutions, as earlier selections are never revisited or refined. This optimisation approach is used to target the joint distribution in Joint Kernel Herding, and the conditional distribution in Average Conditional Kernel Herding.
Gradient Flow [2] methods address the problem of distribution compression by solving a discretised Wasserstein gradient flow of the MMD. In practice, this corresponds to performing gradient descent on all points in the compressed set simultaneously, i.e., solving
$\underset{\bm{Z}\subset\mathbb{R}^{d}}{\arg\min}\;\;\frac{1}{m^{2}}\sum_{i,j=1}^{m}k(\bm{z}_{i},\bm{z}_{j})-\frac{2}{m}\sum_{i=1}^{m}\mathbb{E}_{\bm{x}\sim\mathbb{P}_{X}}\left[k(\bm{z}_{i},\bm{x})\right].$
Under certain technical convexity assumptions on the objective [2], this method can be shown to converge to the global optimum; in practice, however, one should only expect to obtain a locally optimal solution. We take this optimisation approach to target the joint distribution in Joint Kernel Inducing Points, and the conditional distribution in Average Conditional Kernel Inducing Points.
Kernel Thinning [18, 19, 67] constructs a compressed set that is a proper subset of the original dataset, i.e., all points in the compressed set are drawn directly from the input data. This restriction stems from the method’s original motivation of thinning the output of Markov Chain Monte Carlo (MCMC) methods, where subsampling is commonly referred to as standard thinning. Kernel thinning proceeds via a two-stage procedure targeting the MMD. First, the input dataset is probabilistically split into $2^{m}$ MMD-balanced candidate sets, each of size $\lfloor n/2^{m}\rfloor$, discarding the remaining $n-2^{m}\lfloor n/2^{m}\rfloor$ points if $n$ is not evenly divisible. Second, the best candidate set from this partitioning is selected, and then greedily refined by iteratively replacing points with others from the original dataset whenever doing so improves the MMD. This construction enables the derivation of convergence rates that are state of the art in the literature, however practically it has the significant disadvantage of throwing away potentially significant amounts of information in the $n-2^{m}\lfloor n/2^{m}\rfloor$ points.
@@
where $\bm{a},\bm{a}^{\prime}\sim\mathbb{P}_{X}$ and $\bm{b},\bm{b}^{\prime}\sim\mathbb{Q}_{X}$. Much like the MMD, the energy distance is zero if and only if $\mathbb{P}_{X}=\mathbb{Q}_{X}$ [50, Theorem 1]. Although Support Points is not initially expressed in kernel form, the energy distance is in fact equivalent to the MMD for a particular choice of negative-definite kernel [66].